The broad band tremor, which had apparently ceased at the end of the last review period, has started again. An episode of low to moderate tremor commenced at about 07:00 and continued at this level for about 6 hours until approximately 13:00 on 16 February. Small hybrid earthquakes (dominant frequencies of 3-5 Hertz) continue to occur but are only being recorded on the stations nearest to the volcano. Other seismic signals observed throughout the review period were probably the result of small rockfalls. Larger signals of this type were recorded at 00:14, 01:08 and 05:43 on 16 February. The Montserrat seismic network also recorded a magnitude 3.8 regional earthquake at 18:42 on 15 February with a location approximately 240 km south-east of the Island.
An attempt was made to occupy the eastern triangle to Castle Peak today; however, low cloud cover prevented any measurements from being made.
Good views of English's Crater were obtained from the helicopter at 18:00 on the 15 February. Vigorous steam emissions were observed from several parts of the top central parts of the new dome. The south-eastern part of the dome appears to be growing both horizontally and vertically and is now a miniature replica of the southern horizontal protrusion which grew rapidly towards Galways three weeks ago. The southern part, which has stopped growing continues to fall apart along a large fissure which was first noted on 11 February. Several areas of incandescent material was observed within the south-east, west and top central parts of the dome. Rockfalls from the south-eastern part of the dome have begun to extend into the upper parts of the Tar River valley. Low cloud cover and heavy mist prevented any views of English's Crater on 16 February.
Dr Simon Young of the British Geological Survey (BGS) rejoined the monitoring team yesterday evening after an absence of about two weeks. Also arriving yesterday evening were Dr Alan Smith of the University of Puerto Rico and Dr John Roobol of the University of Jeddah. Drs Smith and Roobol will be carrying out investigations of geological deposits on the eastern flanks of the Soufriere Hills volcano. Leaving the island today was Mr Stephane Sol of the SRU after a period of four weeks with the MVO. Mr Sol assisted with the processing of seismic events and with the EDM measurements.